### v4

@@ Abstract
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video can be found at https://youtu.be/y83yw1E-oUo.
@@ 4.2 Evaluation of Mapping and Localization
  Methods Metrics Avg.
Room0
Room1
Room2
Office0
Office1
Office2
Office3
Office4
[…]
24.42
22.12
22.47
24.52
29.07
30.34
19.66
22.23
24.94
NICE-SLAM
[…]
0.809
0.689
0.757
0.814
0.874
0.886
0.797
0.801
0.856
[…]
0.233
0.330
0.271
0.208
0.229
0.181
0.235
0.209
0.198
[…]
30.24
27.27
28.45
29.06
34.14
34.87
28.43
28.76
30.91
Co-SLAM
[…]
0.939
0.910
0.909
0.932
0.961
0.969
0.938
0.941
0.955
[…]
0.252
0.324
0.294
0.266
0.209
0.196
0.258
0.229
0.236
[…]
29.08
25.32
27.77
29.08
33.71
30.20
28.09
28.77
29.71
ESLAM
[…]
0.929
0.875
0.902
0.932
0.960
0.923
0.943
0.948
0.945
[…]
0.336
0.313
0.298
0.248
0.184
0.228
0.241
0.196
0.204
[…]
33.98
32.48
33.72
34.96
38.34
39.04
31.90
29.70
31.68
SplaTAM
[…]
0.969
0.975
0.970
0.982
0.982
0.982
0.965
0.950
0.946
[…]
0.099
0.072
0.096
0.074
0.083
0.093
0.100
0.118
0.155
[…]
34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75
Ours
[…]
0.973 0.976 0.978
0.981
0.984
0.980
0.967 0.966 0.949
[…]
0.096 0.070 0.094 0.070
0.086
0.087
0.101
0.115 0.148
  Methods
Depth L1 [cm]$\downarrow$
ATE Mean [cm]$\downarrow$
ATE RMSE [cm]$\downarrow$
@@ Methods
iMAP
4.645
3.118
4.153
9.92
2.23
1.82
NICE-SLAM
1.903
1.795
2.503
13.70
0.20
0.20
Co-SLAM
1.513
0.935
1.059
17.24
10.20 6.41
ESLAM
1.180
0.520
0.630
18.11
3.62
3.02
Ours 0.356 0.327 0.412
5.27
3.52
2.11
@@ 4.3 Evaluation of Semantic Segmentation
  Methods Avg. mIoUnormal-$\uparrow$
Room0 [%$\uparrow$
Room1 [%$\uparrow$
Room2 [%$\uparrow$
Office0 [%$\uparrow$
NIDS-SLAM
82.37
82.45
84.08
76.99
85.94
DNS-SLAM
84.77
88.32
84.90
81.20
84.66
SNI-SLAM
87.41
88.42
87.43
86.16
87.63

### v5

@@ Abstract
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video is available at https://youtu.be/y83yw1E-oUo.
@@ 4.2 Evaluation of Mapping and Localization
  Methods Metrics Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
@@
[…]
iMAP 4.645 3.118 4.153 9.92 2.23 1.82
NICE-SLAM 1.903 1.795 2.503 13.70 0.20 0.20
Co-SLAM 1.513 0.935 1.059 17.24 10.20 6.41
ESLAM 1.180 0.520 0.630 18.11 3.62 3.02
Ours 0.356 0.327 0.412 5.27 3.52 2.11
@@ 4.3 Evaluation of Semantic Segmentation
[…]
NIDS-SLAM 82.37 82.45 84.08 76.99 85.94
DNS-SLAM 84.77 88.32 84.90 81.20 84.66
SNI-SLAM 87.41 88.42 87.43 86.16 87.63
@@ 4.6 Ablation Study
4.6.1 Effect of Multi-channel Optimization
4.6.2 Effect of Keyframe Optimization